# EP240414a: A Gamma-Ray Burst Jet Weakened by an Extended Circumstellar Material

### v1

@@ 5 Discussion & Conclusions
As shown in Figure 2, our model successfully reproduces the observed multi-wavelength light curves. We found that thermal cooling emission from the non-relativistic outer-cocoon dominates in phase I ($<1$ day), while non-thermal afterglow emission from the mildly relativistic jet-cocoon explains the peak of AT 2024gsa (phase II; $1-10$ days). This is consistent with spectroscopic analyses indicating early thermal emission (in phase I; see van Dalen et al. 2024) and a non-thermal origin for the FBOT peak (phase II; see Sun et al. 2024; van Dalen et al. 2024; Srivastav et al. 2025).
Interestingly, we find that the early cooling emission ($<1$ day) is similar to the first peak in GRB 060218, except that here it blends with the jet afterglow, whereas in GRB 060218, the jet afterglow was too dim to interfere with it (Campana et al., 2006; Nakar, 2015).
@@
From the dynamics of relativistic jets (Matzner 2003; Bromberg et al. 2011b; Nakar 2015; Hamidani et al. 2025), it is understood that CSM environments have observationally measurable consequences on collapsar jets and their transients, allowing us to infer CSM properties and the intrinsic energetics of GRBs. Our study underscores the role of different jet components, ranging from mildly relativistic jets to non-relativistic cocoons; in shaping EP240414a and other GRB transients. We predict that more LGRBs that exhibit CSM interaction signatures will be discovered, particularly through blue thermal cocoon cooling emission on timescales of days, with future high cadence surveys (e.g., LSST, ULTRAST). A notable example is EP250108a, identified at the time of writing (Li et al. 2025), which appears consistent with an FBOT powered by the cooling emission of a completely failed jet (Zhu et al. 2025). Additionally, our results highlight how afterglow emission from a mildly relativistic LGRB jet can mimic (and may account for some) luminous FBOTs. Early multi-wavelength photometric and spectroscopic follow-up of GRBs will refine our understanding of jet-driven components and provide deeper insights into the diversity of their progenitors.
@@ 6 Data availability
The data underlying this article will be shared on reasonable request to the corresponding author.
We thank Katsuaki Asano, Nanae Domoto, Sho Fujibayashi, Kenta Hotokezaka, Susumu Inoue, Christopher Irwin, Wataru Ishizaki, Tomohisa Kawashima, Shota Kisaka, Takashi Nagao, Tatsuya Matsumoto, Ehud Nakar, Ryo Sawada, Jiro Shimoda, Akihiro Suzuki, Kenji Toma, Seiji Toshikage, and Ryo Yamazaki for their fruitful discussions and comments. This research was supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research No. 23K19059, 24K00668, 23H04899 (K. K.), No. 22KJ2643 (Y.S.), and No. 22H00130, 23H05430, 23H04900, 23H01172 (K.I.). Numerical computations were achieved thanks to the following: Cray XC50 of the Center for Computational Astrophysics at the National Astronomical Observatory of Japan, and Cray XC40 at the Yukawa Institute Computer Facility.
@@ Appendix A Extinction correction
Spectroscopy and photometry indicate that SN 2024gsa is similar to SN 1998bw (Sun et al. 2024; van Dalen et al. 2024; although fainter). We estimated the color evolution of SN 2024gsa (in its rest-frame, taking into account the redshift effect and the Galactic extinction $E(B-V)=0.032$; Sun et al. 2024; Dong et al. 2024; Srivastav et al. 2025). We compared its color evolution to that of SN1998bw (Clocchiatti et al. 2011), and to the population of BL Ic SNe (in Taddia et al. 2019). Comparison indicates that SN 2024gsa (as observed) has a much redder rest-frame color. Comparison with the population of BL Ic SNe in Taddia et al. (2019) shows similar $g^{\prime}-r^{\prime}$ color evolution around the peak, but with a larger offset of $\sim 0.6-0.8$ mag toward redder colors. We deduce that this could be due to moderate extinction at the host galaxy (see Prentice et al. 2016; Stritzinger et al. 2018; Taddia et al. 2019).
@@
We conservatively adopt a host extinction of $E(B-V)\sim 0.15$, as a higher value would require more extreme energetics for AT 2024gsa/SN 2024gsa. It should be noted that even such moderate extinction combined with the redshift $z=0.401$ significantly affects the blackbody color because the observed optical emission is emitted in the UV range in the rest-frame, which has a strong dependence on the extinction model ($R_{\rm U}\sim 5$; Schlegel et al. 1998).

### v3

@@ 5 Discussion & Conclusions
Figure 3: Spectral energy distribution as a function of wavelength, in the optical range, showing the contributions from the jet afterglow (dashed), cocoon afterglow (dash-dotted), cocoon cooling emission (dotted), and the sum of all contributions (shaded). The spectral energy distribution is shown at 0.62 days during phase I (first blue peak; blue) and at 4.84 days during phase II (second red peak; red). No correction for Galactic or host-galaxy extinction has been applied. The spectral energy distribution is scaled to the observed spectra (reported in Figure 2 of van Dalen et al. 2025) and shows good agreement.
@@
As shown in Figure 2, our model successfully reproduces the observed multi-wavelength light curves. We found that thermal cooling emission from the non-relativistic outer-cocoon dominates in phase I ($<1$ day), while non-thermal afterglow emission from the mildly relativistic jet-cocoon explains the peak of AT 2024gsa (phase II; $1-10$ days). This is consistent with spectroscopic analyses indicating early thermal emission (in phase I; see van Dalen et al. 2025) and a non-thermal origin for the FBOT peak (phase II; see Sun et al. 2024; van Dalen et al. 2025; Srivastav et al. 2025).
In Figure 3, we show the spectral energy distribution of our model, including contributions from the afterglow (jet and cocoon; non-thermal) and cocoon cooling emission (thermal), at the first peak (0.62 days; phase I) and the second peak (4.84 days; phase II). Comparison with the observed spectra from Figure 2 of van Dalen et al. (2025) shows good agreement between our model and observations at both epochs. [^3] The spectral energy distribution shown in Figure 3 has not been corrected for Galactic or host-galaxy extinction (see Appendix A). Accounting for extinction would slightly reduce the flux and further improve consistency with the observations. This supports the validity of our model and, in particular, demonstrates that the rapid color evolution from blue to red is driven by the short-lived cocoon cooling emission, which dominates only during phase I.
We adopt an on-axis viewing angle in our model ($\theta_{\rm v}\sim 0^{\circ}$), motivated by the very bright optical emission (in phases I and II) and its likely afterglow origin (see Figure 3, and Figure 2 in van Dalen et al. 2025). Our scenario is particularly appealing as the required kinetic energy of the jet in the afterglow phase is of similar order to the energy of the prompt emission ($\sim 10^{50}$ erg; Sun et al. 2024). A slightly off-axis viewing angle remains possible, provided that the observer’s viewing angle is not significantly larger than the jet-cocoon opening angle ($\theta_{\rm v}\lesssim\theta_{\rm j}$). A significantly off-axis configuration ($\theta_{\rm v}\gg\theta_{\rm j}$) would be difficult to reconcile with the observations in our scenario, as it makes the prompt emission and the early afterglow emission (in optical) much dimmer. Different configurations with off-axis jets may also be possible (see Zheng et al. 2025).
@@
Comparing EP240414a to LLGRBs gives a distinct difference in the form of a strong afterglow component that powers the second optical peak (phase II) (Sun et al., 2024; van Dalen et al., 2025; Srivastav et al., 2025). However, if EP240414a were viewed off-axis, at a large viewing angle ($\theta_{v}\gg 1/\Gamma$, with $\Gamma\sim 5$ here), the afterglow-powered optical component (phase II) would be beamed out, and the light curve would evolve from three peaks (phases I to III) into two thermal peaks (phases I and III, originating from cocoon cooling and the supernova, respectively), which would appear as similar to GRB 060218 (Campana et al., 2006; Nakar, 2015).
Given the apparent observational similarity between a completely failed jet (as in GRB 060218) and a barely failed jet viewed off-axis (as in our scenario for EP240414a), despite their fundamentally different physical origins, the question is: how can one distinguish between the two scenarios? A promising diagnostic is radio observation. In our scenario for EP240414a, the off-axis jet-cocoon radio afterglow is expected to exhibit a rising phase followed by a decline at late times (after 10similar-toabsent10$\sim$ 10∼ 10100100100100 days, as determined by the transition time to the mildly-relativistic quasi-spherical state). In contrast, for a completely failed jet such as in GRB 060218, the radio afterglow is expected to be powered by mildly-relativistic, quasi-spherical ejecta, resulting in a steadily declining radio emission (see Figure 2 in Soderberg et al. 2006). Hence, while the optical light curve of an off-axis EP240414a may resemble that of GRB 060218, the radio afterglow provides a key diagnostic to distinguish between a barely failed jet viewed off-axis and a fully failed jet. This highlights the importance of late-time radio observations.
Another EP event of interest is EP250108a, which appears broadly consistent with GRB 060218 (Eyles-Ferris et al., 2025; Rastinejad et al., 2025; Srinivasaragavan et al., 2025; Li et al., 2025), although no radio counterpart has yet been detected. Much later-time radio observations will be required to constrain the jet-cocoon component (see Figure 10 in Eyles-Ferris et al. 2025; Figure 7 in Srinivasaragavan et al. 2025; and Extended Figure 4 in Li et al. 2025).
@@
From the dynamics of relativistic jets (Matzner 2003; Bromberg et al. 2011b; Nakar 2015; Hamidani et al. 2025), it is understood that CSM environments have observationally measurable consequences on collapsar jets and their transients, allowing us to infer CSM properties and the intrinsic energetics of GRBs. Our study underscores the role of different jet components, ranging from mildly relativistic jets to non-relativistic cocoons; in shaping EP240414a and other GRB transients. We predict that more LGRBs that exhibit CSM interaction signatures will be discovered, particularly through blue thermal cocoon cooling emission on timescales of days, with future high cadence surveys (e.g., LSST, ULTRAST). A notable example is EP250108a, identified at the time of writing, which appears consistent with an FBOT powered by the cooling emission of a completely failed jet (Eyles-Ferris et al. 2025; Rastinejad et al. 2025; Srinivasaragavan et al. 2025; Li et al. 2025). Additionally, our results highlight how afterglow emission from a mildly relativistic LGRB jet can mimic (and may account for some) luminous FBOTs. Early multi-wavelength photometric and spectroscopic follow-up of GRBs will refine our understanding of jet-driven components and provide deeper insights into the diversity of their progenitors.
@@ 6 Data Availability
The code used in this article is publicly available via Zenodo: https://doi.org/10.5281/zenodo.15349354 (catalog 10.5281/zenodo.15349354). Additional data underlying this article are available from the corresponding author upon reasonable request.
We thank Katsuaki Asano, Nanae Domoto, Sho Fujibayashi, Kenta Hotokezaka, Susumu Inoue, Christopher Irwin, Wataru Ishizaki, Tomohisa Kawashima, Shota Kisaka, Tatsuya Matsumoto, Takashi Nagao, Ehud Nakar, Hirofumi Noda, Ryo Sawada, Jiro Shimoda, Akihiro Suzuki, Kenji Toma, Seiji Toshikage, and Ryo Yamazaki for their fruitful discussions and comments. This research was supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research No. 23K19059, 24K00668, 23H04899 (K. K.), No. 25KJ0010 (Y.S.), and No. 22H00130, 23H05430, 23H04900, 23H01172 (K.I.). Numerical computations were achieved thanks to the following: Cray XC50 of the Center for Computational Astrophysics at the National Astronomical Observatory of Japan, and Cray XC40 at the Yukawa Institute Computer Facility.
@@ Appendix A Extinction correction
Spectroscopy and photometry indicate that SN 2024gsa is similar to SN 1998bw (Sun et al. 2024; van Dalen et al. 2025; although fainter). We estimated the color evolution of SN 2024gsa (in its rest-frame, taking into account the redshift effect and the Galactic extinction $E(B-V)=0.032$; Sun et al. 2024; Dong et al. 2024; Srivastav et al. 2025). We compared its color evolution to that of SN1998bw (Clocchiatti et al. 2011), and to the population of BL Ic SNe (in Taddia et al. 2019). Comparison indicates that SN 2024gsa (as observed) has a much redder rest-frame color. Comparison with the population of BL Ic SNe in Taddia et al. (2019) shows similar $g^{\prime}-r^{\prime}$ color evolution around the peak, but with a larger offset of $\sim 0.6-0.8$ mag toward redder colors. We deduce that this could be due to moderate extinction at the host galaxy (see Prentice et al. 2016; Stritzinger et al. 2018; Taddia et al. 2019; also see Srinivasaragavan et al. 2024).
@@
We conservatively adopt a host extinction of $E(B-V)\sim 0.15$, as a higher value would require more extreme energetics for AT 2024gsa/SN 2024gsa. We adopt Cardelli et al. (1989) extinction law for both Galactic and host extinction. It should be noted that even such moderate extinction combined with the redshift $z=0.401$ significantly affects the blackbody color because the observed optical emission is emitted in the UV range in the rest-frame, which has a strong dependence on the extinction model ($R_{\rm U}\sim 5$; Schlegel et al. 1998).
Appendix B Summary of Parameters
Table 1: Emission models and their parameters.
Inner- Outer-
Parameters Jet Cocoon Cocoon
Emission Afterglow Afterglow Cooling
Viewing angle [$\theta_{\rm v}$, rad] On-axis ($=0$)
Energy [erg] $1.5\times 10^{50}$ $10^{50}$ $5\times 10^{51}$
Velocity [$\Gamma\beta$] 20 5 $0.4-0.6$
Opening angle [rad] 0.1 0.2 $\pi/2$
Opacity $\kappa$ [cm2 g-1] 0.07
ISM density $n_{0}$ [cm-3] 0.6 0.6
Spectral index $p$ 2.4 2.4
Electron fraction $f_{e}$ 0.2 0.4
Electron energy fraction $\epsilon_{e}$ $3\times 10^{-2}$ $4\times 10^{-1}$
Magnetic energy fraction $\epsilon_{B}$ $2\times 10^{-4}$ $6\times 10^{-2}$